### v1

# Boosting Adversarial Transferability with Spatial Adversarial Alignment

Abstract
Deep neural networks are vulnerable to adversarial examples that exhibit transferability across various models. Numerous approaches are proposed to enhance the transferability of adversarial examples, including advanced optimization, data augmentation, and model modifications. However, these methods still show limited transferability, particularly in cross-architecture scenarios, such as from CNN to ViT. To achieve high transferability, we propose a technique termed Spatial Adversarial Alignment (SAA), which employs an alignment loss and leverages a witness model to fine-tune the surrogate model. Specifically, SAA consists of two key parts: spatial-aware alignment and adversarial-aware alignment. First, we minimize the divergences of features between the two models in both global and local regions, facilitating spatial alignment. Second, we introduce a self-adversarial strategy that leverages adversarial examples to impose further constraints, aligning features from an adversarial perspective. Through this alignment, the surrogate model is trained to concentrate on the common features extracted by the witness model. This facilitates adversarial attacks on these shared features, thereby yielding perturbations that exhibit enhanced transferability. Extensive experiments on various architectures on ImageNet show that aligned surrogate models based on SAA can provide higher transferable adversarial examples, especially in cross-architecture attacks.
1 Introduction
Deep neural networks (DNNs) have been successfully and extensively deployed across security-sensitive applications, including autonomous driving [46, 51, 50], facial verification [53, 52, 34], and video surveillance [17, 54, 18, 14]. However, DNNs exhibit considerable vulnerability to adversarial examples [13, 27, 12, 3, 4, 6, 5], where imperceptible perturbations are introduced into natural images, leading models to produce incorrect predictions. In real-world applications, DNNs are typically concealed from user access, necessitating adversaries to generate adversarial examples within a black-box setting, where no knowledge of the target model’s parameters or architecture is available. Adversarial transferability plays a crucial role in black-box settings as it allows adversaries to effectively compromise target models by employing adversarial examples generated on surrogate models. In black-box settings, adversarial transferability plays a crucial role, which enables adversaries to leverage adversarial examples crafted on surrogate models to effectively attack target models. Thus, generating highly transferable adversarial examples is instrumental in uncovering and understanding the vulnerabilities within DNNs, drawing substantial attention in recent research.
Cross-model transferability has been extensively studied for CNNs [9, 45, 10]. Highly transferable adversarial examples are usually based on advanced optimization [9, 22, 37] and data augmentation [45, 10, 25]. The principle is to alleviate the overfitting of adversarial examples on surrogate models, determining whether the attack can be successfully transferred to the target models. In addition, some model modification methods [44, 15], such as amplifying the gradient on skip connections (the structure in ResNet [16]), can also improve transferability. However, few works explore adversarial transferability on Vision Transformer (ViT) [11] and the performance of existing work extending CNN to ViT is poor due to significant structural differences. Specifically, ViT flattens the image into a sequence of patch tokens and employs multi-head self-attention to capture global relationships among the patches. In contrast, CNNs typically consist of stacked convolutional layers that learn feature relationships progressively through downsampling. Therefore, [43] first empirically analyzes the structure of ViT and propose PNA and PatchOut [43], but there is still much room for improvement in cross-architecture transferability.
In this paper, we argue that unique structural features are critical to cross-architecture adversarial transferability. Given a dataset, various models tend to exhibit analogous decision boundaries [23], arising from their ability to learn similar features. If we can obtain a surrogate model whose features are similar to those of models with different architectures, then the resulting adversarial perturbation can be transferable across different models. A recent technique known as Model Alignment (MA) [26] employs an alignment loss to minimize prediction divergences between surrogate models and witness models, thereby indirectly facilitating the extraction of features that are similarly represented by the witness model. However, directly applying MA to black-box attacks may lead to the degradation of cross-architecture transferability. The main reasons are: (i) Features are not aligned in space. MA only uses the final prediction of the model, but in fact, the spatial features of ViT and CNN are different. It is difficult to directly constrain the similarity of features only by the final logits. (ii) Features are not aligned from the perspective of adversarial features. In addition to the features of clean images, the features of adversarial examples also have similarities across different models and need to be considered.
To overcome these challenges and enhance transferability, we propose a technique called Spatial Adversarial Alignment (SAA), which utilizes an alignment loss from the perspective of spatial and adversarial features and incorporates a witness model to refine the surrogate model. SAA consists of two key parts: spatial-aware alignment and adversarial-aware alignment. In the spatial-aware alignment, in addition to aligning on the final global features, we also focus on the features of local regions. We make local features of CNNs by position to align ViTs’ embeddings at the same position. In the adversarial-aware alignment, we introduce a self-adversarial strategy, which constructs adversarial examples so that the model can learn the differences between different architectures in adversarial features, thereby enabling the model to further capture more common features. Aligned surrogate models by SAA provide promising adversarial transferability and can be seamlessly integrated with existing transfer attacks. Our contributions can be summarized as follows:
  • We reveal for the first time the importance of spatial and adversarial features for cross-architecture transferability.
  • We propose Spatial Adversarial Alignment (SAA), which leverages a witness model to fine-tune the surrogate model via spatial-aware and adversarial-aware alignment to generate highly transferable adversarial examples.
  • Experiments on 6 CNNs and 4 ViTs show that SAA has state-of-the-art adversarial transferability, especially in cross-model transferability. Compared with MA, on ResNet50, the transferability from CNN to ViT is improved by 25.5-39.1%.
2 Related Work
2.1 Transfer attacks on CNNs
Early transfer attacks are mainly conducted between CNNs, and the most popular methods were advanced optimization [9, 22, 37], data augmentation [45, 10, 25], and model modification [44, 15, 43].
Advanced Optimization. [22] compare adversarial attacks to model training: better optimization methods can obtain models with better generalization, and therefore also generate adversarial examples with higher transferability. FGSM [13] is the earliest gradient-based transfer attack, which was then extended to I-FGSM [20]. The subsequent advance optimization further improves the transferability by introducing momentum [9, 22, 39, 37] and smoothness [31].
Data Augmentation. Data augmentation serves as an effective strategy to prevent model overfitting, achieving state-of-the-art performance in model generalization [48, 8]. Building on this principle, numerous adversarial attacks incorporate various transformations to enhance adversarial transferability, including modifications in size [45], scale [22], mixup [40], and frequency domain [25] adjustments. This integration aims to mitigate the overfitting of adversarial examples to the surrogate model, thereby increasing their effectiveness across different models.
Model Modification. According to certain characteristics of the model, modifying the parameters of the surrogate model or changing the forward or backward propagation can also improve the transferability. Skip Gradient Method (SGM) [44] using more gradients from the skip connections rather than the residual modules, allows one to craft adversarial examples with high transferability. Similarly, Linear Backpropagation (LinBP) [15] and Backward Propagation Attack (BPA) [41] concentrate on non-linear activations by modifying the ReLU derivatives to enhance attack transferability. Model Alignment (MA) [26] promotes alignment of model predictions through an alignment loss relative to a witness model, with the aim of capturing shared features across models. However, MA overlooks spatial and adversarial feature alignment across architectures, limiting its effectiveness. Unlike these methods, our SAA requires no modifications to the forward or backpropagation processes, enabling the efficient generation of highly transferable adversarial examples with minimal training overhead. In contrast, LinBP and BPA, involve altering backpropagation or even full model retraining, incurring significantly higher computational costs.
2.2 Transfer attacks on ViTs
Current transfer attacks for ViTs largely adapt methods developed for CNNs. Pay No Attention (PNA) [43] method extends Skip Gradient Method (SGM) to ViTs by omitting the gradient computation of attention blocks during back-propagation, thereby enhancing adversarial transferability. PatchOut [43] strategy selects a random subset of image patches to compute the gradient at each iteration, functioning as an image transformation technique to increase transferability. Then, Self-Ensemble (SE) [29] approach employs the class token at each layer with a shared classification head to create an ensemble model, facilitating optimized perturbation; however, many ViTs, such as Visformer [2] and CaiT [36], lack sufficient class tokens to build this ensemble. Additionally, Token Refinement (TR) [29] module fine-tunes class tokens to further boost transferability. Recently, Token Gradient Regularization (TGR) [49] works from the perspective of variance reduction, stabilizing the gradient direction to prevent adversarial examples from getting stuck in poor local optima. Distinct from these approaches, SAA is the first method to specifically analyze architectural differences across models. By leveraging shared features between different architectures, SAA enables the creation of more generalized surrogate models that integrate seamlessly with optimization and data augmentation methods, ultimately achieving state-of-the-art transferability.
3 Methodology
3.1 Preliminaries
In this paper, we focus on the image classification task on DNNs. Let $f_{\theta}(\cdot)$ represent a DNN-based classifier with parameters $\theta$. We denote the clean image as $x$ and its corresponding ground-truth label as $y$. Following [9, 45, 10], we evaluate the adversarial transferability under untargeted adversarial attacks with $l_{\infty}$ norm. Therefore, the goal of transfer attacks is to add an adversarial perturbation to the clean image $x$ based on the information of the surrogate model $f_{\theta_{s}}(\cdot)$ to obtain the adversarial example $x_{adv}$, so that $f_{\theta_{s}}(x_{adv})\neq y$ subject to the constraint that $||x_{adv}-x||_{\infty}\leq~{}\epsilon$. In the black-box setting, no information about the target model—such as its architecture, weights, or gradients—is accessible. Therefore, adversarial examples are generated solely by utilizing a surrogate model $f_{\theta_{s}}(\cdot)$, leveraging their transferability to deceive the target model $f_{\theta_{t}}(\cdot)$.
3.2 Spatial Adversarial Alignment (SAA)
Spatial Adversarial Alignment (SAA) employs an alignment loss tailored to both spatial and adversarial feature perspectives, incorporating a witness model to fine-tune the surrogate model. SAA aims to adjust the surrogate model to extract features closely aligned with those of the witness model, capturing both spatial and adversarial features shared across models. As shown in Figure 1, SAA consists of two parts, namely spatial-aware alignment and adversarial-aware alignment.
Spatial-aware Alignment. The purpose of spatial-aware alignment is to make the surrogate and witness models more consistent in the feature space. Naturally, the most intuitive approach to aligning the feature distributions of two models is to minimize the distance between their final outputs. However, when the models exhibit significant architectural differences, ensuring output similarity alone is insufficient to achieve alignment in intermediate features. In black-box attacks, where the details of the target model’s architecture are unknown, this issue becomes more pronounced. Taking the challenging scenario of CNN to ViT as an example, their intermediate layer features differ substantially in semantic levels. This discrepancy arises primarily from differences in receptive fields, stacking methodologies, and normalization techniques between CNNs and ViTs. Therefore, relying solely on output alignment for model fine-tuning indirectly captures some common features, but this approach can, in certain cases, result in degraded transferability, as observed in methods like Model Alignment (MA) [26].
Therefore, in addition to aligning on the final global features, we also need focus on the features of local regions. For ease of understanding, we define the global features $f_{\theta}(x)$ as the logits of the model corresponding to the input $x$. For CNNs, it is the output of features by the last layer. For ViTs, it refers to the final embedding of the [CLS] token after the MLP block. First, we perform alignment at the global feature level by defining an alignment loss between the surrogate model and witness model at the output layer:
$\mathcal{L}_{global}(x;\theta_{s})=D_{\mathrm{KL}}(f_{\theta_{s}}(x),f_{\theta_{w}}(x)),$ (1)
where $D_{\mathrm{KL}}$ measures the feature divergence with Kullback-Leibler (KL) divergence.
Figure 1: Spatial Adversarial Alignment (SAA) consists of two parts: spatial-aware alignment and adversarial-aware alignment. Initially, we aim to minimize the feature divergences between the two models across both global and local regions, thereby promoting spatial alignment. Subsequently, we introduce a self-adversarial strategy that utilizes adversarial examples to impose additional constraints, aligning the adversarial features.
Next, we align the models at the local feature level. Here, we define $z_{\theta}(x)$ as local features. For CNNs, $z_{\theta}(x)^{B\times C\times H\times W}$ is the features of the last layer before passing through the global average pooling layer. For ViTs, $z_{\theta}(x)^{B\times C^{\prime}\times H\times W}$***Generally, ViT’s patch embeddings $z(x)$ is $(B,N,C’)$ ’ ) by default. We first transform it to $(B,C^{\prime},H’,W’)$ ’ ), where $N=H^{\prime}\times W^{\prime}$. Then, we perform an adaptive pooling operation to transform it to $(B,C^{\prime},H,W)$. is the embeddings of the patch tokens after passing through the MLP except for the [CLS] token, where each patch token corresponds to a specific spatial region in the input. Each spatial position $(h,w)$ within this feature map is treated as a distinct local region and the feature for each local region is $z_{\theta}^{[q]}(x)$, where $q=\{1,2,...,H\times W\}$. Let $z_{\theta_{s}}^{[q]}(x)$ and $z_{\theta_{w}}^{[q]}(x)$ denote the local prediction (logits) associated with each local region $q$ for the surrogate model and witness model, respectively. Next, we compute the pseudo-labels of the local region $q$ based on the local features of the witness models. We denote this pseudo-label as $\hat{y}^{[q]}_{\theta_{w}}$, which is obtained by taking the $\arg\max$ over the logits from the witness model: $\hat{y}^{[q]}_{\theta_{w}}=\arg\max(f_{\theta_{w}}(z_{\theta_{w}}^{[q]}(x)))$. Then, we use this pseudo-label to supervise the learning of the local feature of the surrogate model. To achieve local alignment, we minimize the divergence between the logits of corresponding local regions, so the local alignment loss is expressed as:
$\small\mathcal{L}_{local}(x;\theta_{s})=\frac{1}{HW}\sum^{HW}_{q=1}D_{\mathrm{CE}}(z_{\theta_{s}}^{[q]}(x),\arg\max(f_{\theta_{w}}(z_{\theta_{w}}^{[q]}(x)))),$ (2)
where $D_{CE}$ is the cross-entropy loss. Therefore, the spatial-aware alignment loss is expressed as:
$\mathcal{L}_{SA}(x;\theta_{s})=\mathcal{L}_{global}(x;\theta_{s})+\gamma\cdot\mathcal{L}_{local}(x;\theta_{s}),$ (3)
where $\gamma$ is the spatial factor. By minimizing this spatial-aware alignment loss, we encourage the surrogate model to produce features in both global and local regions that are consistent with those of the witness model, even across different architectures.
Figure 2: Grad-CAM visualizations comparing the feature distribution of unaligned and aligned surrogate models (Res50) on clean inputs and adversarial examples (generated by SSA-DI-TI-MI).
Adversarial-aware Alignment. The relationship between features and adversarial vulnerability is highly significant. Some hypotheses [47, 30] propose that adversarial examples possess distinct feature distributions compared to normal examples, which may inherently predispose models to adversarial vulnerability—a notion supported by several studies [1, 25]. Beyond normal examples, learning adversarial features may offer a way to capture shared features between surrogate models and witness models. Furthermore, [10] suggests that models trained with adversarial examples focus on more discriminative regions within images, displaying feature recognition patterns distinct from those of normally trained models. Thus, adversarial examples play a crucial role in achieving model alignment.
In our adversarial-aware alignment, we introduce a self-adversarial strategy that constructs adversarial examples of the surrogate model to enable the model to discern architectural differences in adversarial features effectively. Specifically, we leverage the gradients to iteratively generate adversarial examples under the supervision of the global features of the witness model. Assuming $x_{adv}^{(0)}=x$, we define the adversarial example $x_{adv}^{(t+1)}$ of the surrogate model as:
$\footnotesize x_{adv}^{(t+1)}=\Pi_{x,\epsilon}\left(x_{\text{adv}}^{(t)}+\alpha\cdot\text{sign}\left(\nabla_{x}D_{\text{KL}}\left(f_{\theta_{s}}(x_{\text{adv}}^{(t)}),f_{\theta_{w}}(x)\right)\right)\right),$ (4)
where $D_{\mathrm{KL}}$ denotes the KL divergence, $x_{adv}^{(t)}$ denotes the adversarial example at iteration $t$, $\alpha$ is the step size, and $\Pi_{\epsilon}$ projects the adversarial example onto an $\epsilon$-bounded neighborhood around the original input $x$.
Once the adversarial example $x_{adv}$ is generated, we also perform adversarial-aware alignment on the adversarial examples from global and local features to further align the surrogate and witness models. The loss of the adversarial-aware alignment is expressed as:
$\mathcal{L_{AA}}(x_{adv};\theta_{s})=\mathcal{L}_{global}(x_{adv};\theta_{s})+\omega\cdot\mathcal{L}_{local}(x_{adv};\theta_{s}),$ (5)
where $\omega$ is the adversarial factor.
Optimization. Combining spatial-aware and adversarial-aware alignment, the final optimization goal of spatial-adversarial alignment is:
$\mathcal{L}_{SAA}(x;\theta_{s})=\mathcal{L}_{SA}(x;\theta_{s})+\kappa\cdot\mathcal{L}_{AA}(x_{adv};\theta_{s}),$ (6)
where $\kappa$ is the alignment factor to balance the two alignments. If not otherwise stated, we define $\gamma=0.2$, $\omega=0.02$, and $\kappa=0.02$ in this paper.
Spatial-adversarial alignment facilitates the alignment of the surrogate model with a collection of witness models to further improve the adversarial transferability. This process involves utilizing a set of witness models, denoted as $\Theta$, where the cardinality of the set is represented by $|\Theta|$. Therefore, the parameter update rule for the surrogate model, based on stochastic gradient descent (SGD), can be expressed as follows:
$\theta_{s}^{(t+1)}=\theta_{s}^{(t)}-\eta\cdot\frac{1}{|\mathcal{B}||\Theta|}\sum\limits_{x\in\mathcal{B}}\sum\limits_{\theta_{w}\in\Theta}\nabla_{\theta_{s}^{(t)}}\mathcal{L}_{SAA}(x;\theta_{s}),$ (7)
where $t$ is the epoch, $\eta$ is the learning rate and $\mathcal{B}$ means the mini-batch samples.
3.3 A Close Look at SAA
To verify whether SAA significantly improves the spatial and adversarial features after model alignment, we conduct quantitative and qualitative analyses based on the models before and after alignment. We randomly sample 100 images from ImageNet val and then compute the cosine similarity between the global features of the surrogate models before and after applying SAA with different witness models. Table 1 demonstrates that, whether for clean images or adversarial examples (generated by SSA-DI-TI-MI), the feature similarity improves after alignment. Notably, when the surrogate model is ViT-B, the improvement in similarity is even more pronounced. This result suggests that, after applying SAA, the aligned surrogate models effectively capture features shared with the witness model, providing strong evidence of the alignment’s success.
Table 1: Cosine similarity of global features of surrogate models.
Surrogate Witness Clean Adv
Unaligned Aligned Unaligned Aligned
Res50 Res50 1.0000 0.9949 1.0000 0.9922
DN121 0.0573 0.1153 0.0700 0.1328
ViT-B 0.0533 0.1408 0.0452 0.1191
Swin-B 0.0352 0.0448 0.0369 0.0551
ViT-B Res50 0.0566 0.1323 0.0672 0.1544
DN121 0.4016 0.6278 0.4121 0.6551
ViT-B 1.0000 0.9706 1.0000 0.9728
Swin-B 0.3058 0.5115 0.3169 0.4257
Then, we use Grad-CAM [32]’s heatmaps to simulate the feature distribution of the model, as shown as Figure 2. For the clean inputs (first four cols), the heatmaps generated by the unaligned surrogate model (2-nd col) primarily focus on local regions of the object. In contrast, the aligned surrogate model (4-th col) heatmaps demonstrate more diffuse attention spread across the entire object, similar to that of the witness model (ViT-B, 3-rd col), which shows aligned surrogate models learn the common spatial features. For adversarial examples (last four cols), the 5-th and 7th cols display the heatmaps of adversarial examples generated by the unaligned and aligned surrogate models, respectively. The 6-th and 8-th show the witness model’s responses to these adversarial examples. Notably, the adversarial examples generated by the unaligned surrogate model fail to effectively transfer to the witness model (6-th col) due to still focusing on the target subject, indicating limited cross-model transferability. In contrast, adversarial examples generated by the aligned surrogate model (7-th col) successfully transferred to the witness model (8-th col) as the features are spread out, demonstrating enhanced cross-model transferability achieved through SAA.
4 Experiments
4.1 Experimental Setup
Datasets. Our experiments utilize the ImageNet-compatible dataset [21], a widely adopted subset containing 1,000 images from the ImageNet validation set [7]. This dataset is commonly used in adversarial robustness studies, such as those in [45, 9, 10].
Models. To assess the adversarial transferability of different network architectures, we focus on convolutional neural networks (CNNs) and vision transformers (ViTs) as the target models. For CNNs, we select the typically trained ResNet-18 (Res18), ResNet-50 (Res50) and ResNet-101 (Res101) [16], VGG-19 [33], DenseNet-121 (DN121) [19], and Inception-v3 (Inc-v3) [35]. For ViTs, we evaluate the Vision Transformer (ViT-B) [11], Swin Transformer (Swin-B) [24], Pyramid Vision Transformer (PVT-v2) [38], and MobileViT-s (MobViT) [28].
Metric. Adversarial transferability is quantified by calculating the average attack success rate (Avg. ASR, %) across target models (excluding the surrogate model), with a higher success rate signifying enhanced transferability. In the paper, ’n/a’ is defined as the average attack success rate obtained by generating adversarial examples using the surrogate model without any alignment, serving as a baseline.
Table 2: Comparison of adversarial transferability on different alignment methods.
Surrogate Witness Attack Target Model Avg. ASR (%)
CNNs ViTs
Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT
Res50 n/a MI 57.7 99.9 58.1 54.2 55.1 39.0 9.4 33.0 38.0 35.7 42.2
Res50 MA 60.4 99.8 56.4 60.3 67.3 44.3 12.1 35.6 37.2 39.0 45.8
SAA 77.5 100.0 71.7 72.0 77.0 58.7 19.9 47.8 51.8 52.8 58.8
DN121 MA 83.1 96.7 75.8 82.3 87.1 64.0 19.8 49.6 54.5 59.2 63.9
SAA 92.4 98.6 87.1 90.3 94.6 77.8 30.5 64.2 69.2 76.5 75.8
ViT-B MA 74.2 99.2 63.5 69.3 72.8 51.5 18.5 41.5 42.7 47.4 53.5
SAA 84.1 99.6 74.7 80.3 81.8 65.7 24.3 48.7 52.9 62.5 63.9
Swin-B MA 64.7 90.4 50.6 61.2 61.6 42.3 10.7 33.6 36.4 38.8 44.4
SAA 79.7 95.7 66.4 74.1 75.9 56.9 19.6 41.9 47.9 55.2 57.5
DN121 n/a MI 84.4 69.6 54.8 76.6 100.0 56.5 16.2 42.1 43.9 52.6 58.6
Res50 MA 88.8 65.8 49.8 78.3 100.0 55.7 14.9 40.0 38.2 53.0 57.6
SAA 95.7 80.3 70.9 87.7 99.6 78.2 31.4 55.0 56.6 72.1 71.9
DN121 MA 79.3 69.5 54.9 78.8 100.0 55.4 12.3 39.9 41.7 52.4 57.2
SAA 90.1 81.4 71.5 87.9 99.9 75.0 25.3 54.9 56.4 69.7 70.1
ViT-B MA 89.6 80.6 71.1 88.1 100.0 70.0 22.4 53.9 58.4 68.7 69.1
SAA 94.0 83.6 75.2 90.3 99.8 81.7 27.2 58.0 59.1 80.3 74.0
Swin-B MA 88.3 63.9 49.2 77.8 100.0 54.4 14.5 38.1 37.5 52.2 56.9
SAA 94.8 80.0 69.1 88.0 99.7 74.4 26.3 50.2 51.6 71.2 69.5
ViT-B n/a MI 48.6 41.0 31.9 58.1 50.8 44.4 100.0 58.1 44.7 47.4 53.8
Res50 MA 65.8 51.9 43.0 66.9 60.0 51.5 99.7 66.0 57.3 57.6 63.1
SAA 87.3 80.6 75.9 86.1 87.5 78.0 99.9 91.0 85.4 85.2 86.3
DN121 MA 88.9 73.6 67.6 89.1 87.8 74.9 100.0 84.9 78.3 82.7 83.8
SAA 94.7 86.8 81.3 93.8 94.0 86.6 99.8 91.3 87.2 90.7 91.0
ViT-B MA 54.3 37.2 29.2 56.0 47.9 42.4 100.0 52.8 42.3 47.4 52.5
SAA 62.7 46.9 40.9 64.2 57.5 52.1 100.0 59.5 52.6 58.8 60.9
Swin-B MA 66.6 46.0 38.3 67.5 57.6 50.8 99.7 65.7 53.6 58.2 62.0
SAA 83.0 69.2 65.6 81.2 78.1 74.2 99.3 84.3 76.0 78.3 80.0
Swin-B n/a MI 48.2 31.3 20.0 49.3 34.3 29.7 13.9 100.0 45.4 41.4 42.5
Res50 MA 54.8 50.9 36.5 61.8 49.3 40.1 32.4 100.0 68.5 55.6 55.4
SAA 90.3 85.3 77.9 92.0 88.4 76.4 64.7 99.9 92.4 89.0 85.7
DN121 MA 77.3 71.4 60.0 83.9 76.4 60.7 49.0 100.0 85.9 80.9 74.9
SAA 94.9 91.7 85.8 96.5 94.7 85.6 74.3 100.0 95.5 95.4 91.4
ViT-B MA 62.6 52.8 40.0 66.2 55.3 45.6 34.0 100.0 70.1 61.8 59.5
SAA 84.7 80.0 74.2 90.5 85.5 75.7 70.3 100.0 91.9 88.6 84.6
Swin-B MA 62.6 52.8 40.0 66.2 55.3 45.6 34.0 100.0 70.1 61.8 59.5
SAA 73.5 58.9 45.7 79.2 63.4 50.7 32.8 100.0 75.8 71.4 65.8
Implementation Details. In our experiments, we select the MI [9] attack as the baseline for generating adversarial examples with high transferability, as it is widely recognized within the field of adversarial transferability [10, 45, 39, 25, 40, 37, 22, 42, 44, 43]. For MI, we set the perturbation magnitude $\epsilon=16$, perform 10 iterations, with a step size of $\frac{16}{10}=1.6$, and use a momentum $\mu=1$. During the Spatial Adversarial Alignment, all surrogate models are fine-tuned for one epoch using stochastic gradient descent (SGD) with a momentum of 0.9, and no learning rate adjustments are applied. It is important to note that no additional data is used for fine-tuning, as it relies solely on the same training data used for both the surrogate and witness models. The settings for Model Alignment (MA) [26] are consistent with the parameters specified in the original paper.
Table 3: SAA has stronger adversarial transferability after being compatible with existing transfer attacks.
Target Model
CNNs ViTs
Attack Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT Avg. ASR (%)
MI 57.7 99.9 58.1 54.2 55.1 39.0 26.3 15.1 38.0 35.7 42.1
MI-SAA 84.1 99.6 74.7 80.3 81.8 65.7 53.1 34.8 52.9 62.5 65.5
NI 58.9 100.0 63.2 59.3 61.4 40.0 27.3 15.2 41.8 38.1 45.0
NI-SAA 86.1 99.9 76.3 82.2 83.7 67.6 53.2 35.6 55.7 64.8 67.2
GI 57.3 100.0 62.3 60.5 60.5 40.7 29.5 17.3 40.7 39.6 45.4
GI-SAA 86.5 99.7 78.8 83.9 84.8 70.4 58.8 39.6 55.8 66.3 69.4
DI 44.1 95.8 41.7 56.1 44.2 26.1 16.5 7.7 36.7 35.1 34.2
DI-SAA 74.6 94.4 61.1 81.1 73.2 53.1 34.9 19.9 50.8 63.6 56.9
TI 38.5 99.9 37.8 33.9 36.1 24.2 16.8 7.9 29.0 21.1 27.3
TI-SAA 59.7 94.9 35.2 50.6 54.5 40.6 29.2 13.6 29.3 33.6 38.5
SSA 75.8 99.9 78.6 76.0 77.8 57.0 35.3 24.1 55.0 50.5 58.9
SSA-SAA 91.5 99.5 77.8 85.7 88.4 74.9 54.4 36.9 57.1 66.0 70.3
DI-MI 65.5 97.0 65.0 74.7 65.7 49.1 34.5 22.1 54.9 57.9 54.4
DI-MI-SAA 91.9 98.7 84.9 94.1 90.5 78.3 64.9 46.2 76.1 86.8 79.3
TI-MI 61.4 99.9 60.5 60.9 60.9 44.3 35.5 19.6 42.3 41.8 47.5
TI-MI-SAA 84.8 99.3 71.9 79.0 81.8 69.1 58.9 38.7 52.8 62.4 66.6
SSA-MI 89.6 99.9 92.2 89.5 91.0 77.6 57.8 49.4 76.4 76.3 77.8
SSA-MI-SAA 96.3 99.8 95.6 96.5 97.2 91.5 81.3 68.2 80.4 88.4 88.4
SSA-DI-TI-MI 93.5 98.5 92.3 95.0 93.7 85.5 78.9 66.5 87.1 89.8 86.9
SSA-DI-TI-MI-SAA 97.5 98.8 93.8 97.6 96.7 94.3 90.6 78.0 84.4 94.7 92.0
Table 4: SAA further improves the adversarial transferability of adversarial attacks on ViTs.
Attack Target Model Avg. ASR (%)
CNNs ViTs
Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT
SGM 82.9 67.6 59.4 81.2 75.4 71.3 99.7 83.3 72.7 78.8 78.3
SGM-SAA 91.1 79.8 73.3 87.5 87.3 80.9 99.5 90.5 82.6 86.3 86.6
PatchOut 45.6 27.4 20.3 45.5 36.1 33.9 93.0 41.0 34.2 40.5 43.3
PatchOut-SAA 76.5 72.4 70.3 79.4 78.1 71.3 94.7 83.6 77.2 76.8 78.7
PNA 61.2 45.0 38.1 60.8 54.8 49.0 99.6 66.3 55.8 56.8 60.3
PNA-SAA 82.7 78.1 73.4 85.6 84.0 75.1 97.4 89.3 80.3 82.1 83.3
TGR 74.0 55.6 48.4 73.2 66.6 59.0 99.3 74.5 61.6 69.6 69.6
TGR-SAA 85.9 78.1 71.5 87.4 85.6 79.6 99.3 89.1 81.0 86.2 85.1
4.2 Performance Comparison
Performance comparison with alignment methods. We first compare with existing alignment methods [26], where adversarial examples are generated based on MI [9]. Table 2 illustrates the performance difference between MA and SAA in terms of adversarial transferability, with SAA demonstrating a significant advantage over MA. For instance, when the surrogate model is Res50, and the witness model is also Res50, SAA achieves a 16.6% improvement in average ASR over the original surrogate model, compared to a modest 3.6% improvement with MA. This highlights that SAA, even without introducing additional information, enhances adversarial transferability through the alignment of adversarial features. Furthermore, when the witness models are DN121, ViT-B, and Swin-B, SAA outperforms MA by 11.9%, 10.4%, and 13.1%, respectively. In addition to the remarkable adversarial transferability that SAA provides, we make two other key observations: (i) MA only considers global features, which makes it difficult to align features between models with large differences, which may lead to a decrease in transferability. When the surrogate model is DN121 and the witness model is Swin-B, the ASR of Vit-B, Swin-B, PVT-v2, and MobViT is not as good as the origin DN121, which shows that relying solely on global features for alignment is not enough, and can only achieve poor transferability. (ii) SAA provides strong cross-architecture transferability. When the surrogate model is Res50, and the witness models are Res50, DN121, ViT-B, and Swin-B, the transferability of SAA on ViTs is improved by 39.06%, 31.29%, 25.51%, and 37.74% respectively compared with MA itself, and it also has high transferability between CNNs.
Table 5: Ablation study on alignment modules.
Module Target Model Avg. ASR (%)
Spatial Adversarial CNNs ViTs
Global Local Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT
57.7 99.9 58.1 54.2 55.1 39.0 9.4 33.0 38.0 35.7 42.2
60.4 99.8 56.4 60.3 67.3 44.3 12.1 35.6 37.2 39.0 45.8
67.5 98.0 58.1 69.1 70.1 49.9 12.2 37.7 38.9 47.7 50.1
81.6 97.9 68.1 53.0 79.8 63.6 25.5 47.9 48.7 57.3 58.4
84.1 99.6 74.7 80.3 81.8 65.7 24.3 48.7 52.9 62.5 63.9
Performance comparison with transfer attacks. Aligned surrogate models by SAA have great potential for adversarial transferability, so existing transfer attacks such as advanced optimization and data augmentation can further improve transferability. Here, we choose Res50 as the surrogate model and ViT-B as the witness model, and superimpose them with MI [9], NI [22], GI [37], DI [45], TI [10] and SSA [25] to evaluate the transferability, as shown in Table 3. Taking GI and SSA as examples, the transferability of the model after SAA is improved by 24.0% and 11.4% respectively compared with the origin surrogate model, which is a very significant improvement. When multiple attacks are integrated, such as SSA-DI-TI-MI, SAA further enhances the transferability by 5.1%, achieving an impressive 92.0% ASR, which closely approaches the performance of white-box attacks. This indicates that SAA substantially narrows the performance gap between black-box and white-box attacks, thereby facilitating a more comprehensive evaluation of the adversarial robustness of existing models.
Performance comparison with attacks on ViTs. To further explore the cross-architecture transferability, we evaluate adversarial attacks on ViTs, including SGM [44], PatchOut [43], PNA [43], and TGR [49]. Here, we choose the surrogate model as ViT-B and the witness model as Res50. In PatchOut, SAA first improves the transferability between ViTs, for example, the ASR from ViT-B to Swin-B is improved by 42.6%. Secondly, SAA greatly improves the transferability from ViT to CNN, for example, the ASR is improved by 45.0% and 42.0% on Res50 and DN121 respectively. On SGM, PNA, and TGR, SAA also achieves stronger cross-architecture transferability without modifying the forward and backward propagation of the model.
4.3 Ablation Studies
We select ResNet-50 as the surrogate model and ViT-B as the witness model for ablation studies.
Alignment Module. In spatial-aware alignment, ‘global’ represents $\mathcal{L}_{global}$ (Equation 1), while ‘local’ represents $\mathcal{L}_{local}$ (Equation 2). Similarly, ‘adversarial’ represents $\mathcal{L}_{AA}$ (Eqution 5) of adversarial-aware alignment. As shown in Table 5, when global features are introduced into the alignment, the transferability of the aligned surrogate model will increase by 3.6%. Based on $\mathcal{L}_{global}$, when only local features are introduced, the overall transferability is improved by 4.3% due to better alignment of features of different architectures in local regions, especially by 8.7% on MobViT. Based on $\mathcal{L}_{global}$, when only adversarial features are introduced, the transferability is greatly improved, reaching 12.6% ASR, and the improvement is significant on ViTs. Finally, by integrating all features, the aligned surrogate model achieves state-of-the-art transferability. Through the above experiments, we illustrate the effectiveness of spatial-aware and adversarial-aware alignment.
Training Epochs. In Section 4.2, we reveal the powerful potential of SAA for adversarial transferability after training for only one epoch. Furthermore, we explore the performance difference after training for multiple epochs, as shown in Figure 3. We calculate the average attack success rate except for the Res50 surrogate model itself and find that with the increase of epochs, the adversarial transferability of the aligned surrogate model is further improved, reaching convergence around the 9-th epoch. Compared with MA, SAA can achieve higher transferability in small epochs, and after multiple rounds of training, the transferability has a higher upper limit, which shows the importance of using spatial and adversarial features for model alignment.
Figure 3: Ablation study on training epochs.
5 Conclusions
In this study, we introduce a novel technique called Spatial Adversarial Alignment (SAA), which incorporates an alignment loss function and utilizes a witness model to fine-tune a surrogate model by focusing on both spatial-aware and adversarial-aware alignments. Through comprehensive experimental analysis, we demonstrate that leveraging these spatial and adversarial features for model alignment significantly enhances the adversarial transferability of surrogate models, with a particularly pronounced improvement in their cross-architecture capabilities. The proposed SAA method not only integrates seamlessly with existing transfer attack strategies but also further amplifies adversarial transferability, thereby contributing to a more complete evaluation of the adversarial robustness of current DNNs.
Boarder Impacts. The adversarial examples generated by surrogate models after the application of SAA exhibit enhanced adversarial transferability, with particularly improved cross-architecture capabilities. This augmented transferability poses a significant threat to the deployment of deep learning models in real-world applications. Additionally, through empirical experiments, we demonstrate the critical role of spatial and adversarial features in improving transferability. While a comprehensive theoretical analysis is still lacking, these findings highlight the need for further investigation into adversarial robustness and the development of effective defense mechanisms.